AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 26 of 33  ELOG logo
New entries since:Thu Jan 1 01:00:00 1970
ID Date Author Subject Text Attachments
  156   Wed Nov 17 07:12:22 2021 RSS CB + ML (remotely)CARME baking with tape continues 17 NovCARME baking with heating tape continues

Left overnight at -

Ch1-3: 140 degrees
 7x 
  155   Tue Nov 16 12:05:13 2021 CB RSSCARME baking with tape continuesCARME baking with heating tape continues

Left overnight at 100C. Turned
up this morning at 9:00 to 120C.
 CARMEtemps.png20211116_132048.jpg20211116_155138.jpg20211116_155857.jpg 
  154   Fri Nov 12 08:54:11 2021 RSS ML + CB (remotely) NK AK NPFirst CARME baking attempt07:50 Small leaks (near the motors and
upper valve) were detected in the tent which
were covered with Al foil.
 Capture.PNG200W_heaters.svg40W_heater.svg 
  153   Thu Nov 11 14:01:00 2021 AK, MLHeater channel layoutsSchematic layout of heaters and channel
assignment attached. Table kindly provided
by AK. I hope the conversion from excel to
 2020.11.05_Cryring-YR09-baking-layout.pdf 
  151   Thu Nov 11 11:44:58 2021 RSS + ML OG + CB (remotely)CARME baking startedWith the help of OG, the baking tent was
built. MG started baking around 11:30 am
with a rate of 0.2 K/minute.
 IMG_5535.JPGIMG_5537.JPGIMG_5531.JPGIMG_5534.JPGIMG_5539.JPG 
  150   Wed Nov 10 18:38:16 2021 RSS TD + ML OG AK NKCARME wrapped in heating tapeHeating and electronic experts were in
CRYRING and were unhappy with the damaged
wire. Heaters were opened by OG, but no damage
 IMG_5514(1).JPGIMG_5517(1).JPGIMG_5518.JPGIMG_5520.JPGIMG_5521(1).JPG 
  149   Wed Nov 10 18:24:12 2021 RSSThermocouple labelling1. Detector

2. Chamber back

3. Strain relief
 labeling.JPG.svgLabelling.svgIMG_5512.JPG 
  147   Wed Nov 10 14:41:12 2021 TD RSS + CB (remotely)Wrapping CARME in heating tapeML reports that baking of the upstream section
(gas jet interaction chamber plus e-target)
has commenced - initially to 50 deg C to
  
  146   Tue Nov 9 15:03:29 2021 RSS + OGBurnt cables issuesUpper part of the frame was mounted. In
order to do so, the valve was closed, turbo
wire and hose were disconnected and then
 7x 
  145   Mon Nov 8 17:19:02 2021 RSS + OGNeg and ion cables connectedNeg and ion pump cables were labeled and
connected.

Kapton tape wrapped around the
 IMG_5480.JPGIMG_5483.JPGIMG_5488.JPG 
  144   Sat Nov 6 14:21:10 2021 CB RSSPreparation for bakeout tent mounting + to do for next weekConnected pneumatic hoses prepared by Gleb.
Opened flow to pressure reducer set at 2
bar. Ok. Attempted to operate pneumatic motors.
  
  143   Fri Nov 5 12:27:11 2021 CB RSS + DR NP ML USFirst pump down of YR09Tightened all flanges.

Added Pirani gauge and third IE514
to interaction chamber.
 20211105_131748.jpg20211105_135918.jpg20211105_135929.jpg20211105_140451.jpg20211105_150616.jpg 
  142   Wed Nov 3 16:39:01 2021 CB RSS + NP DR NK MLInteraction chamber lifted in placePrepared to move interaction chamber in
place. Ok CARME side, flange has M10 bolts.
On interaction chamber side i.e. bellows
 20211103_161553.jpg20211103_161813.jpg20211103_175457.jpg 
  141   Wed Nov 3 11:18:11 2021 CB RSS + DR NK MLLeak closedFound CARME at 1.0E-6 mbar. Beginning leak
test.

 
 20211103_110244.jpg20211103_110414.jpg 
  140   Tue Nov 2 12:45:36 2021 CB + DR NKIssues with CARME feetAttempted to close CARME as carefully as
possible.

Swapped out one foot for another
 20211102_134446.jpg 
  139   Mon Nov 1 14:48:16 2021 CB + ML DRCARME openedConfirmed leak as per previous Elog.

Tightened to 100 Nm - no effects.

Tightened to 130 Nm - bottom leak
  
  138   Sun Oct 31 13:55:13 2021 CB OHLeak testsFound CARME at 2E-7.

Closed valve on top of roughing
pump. Clamp was not tight. Pressure still
 20211031_155820.jpg 
  137   Sun Oct 31 08:05:15 2021 CB OH + DR MLCARME closed again + feedthroughs labeled(Fri 29 October)

Cleaned wire seal recess and labeled
thermocouples and flanges to identify p-n
 11x 
  136   Sun Oct 31 07:55:39 2021 CB JM + OHStrain relief mounted(Thu 28 October)

All strain relief extension plates
mounted behind the wires, using SS countersunk
 20211028_190026.jpg20211028_192251.jpg20211028_192210.jpg 
  135   Sun Oct 31 07:41:20 2021 CB JM + RSSWires plugged in DSSDs (Wed 27 October)

Started with top-most flange. Attemptedto
plug in MACOR lead in DSSD plug in preparation
 8x 
ELOG V3.1.4-unknown